Pick out the solution from the values given in the bracket next to each equation. Show that the other values do not satisfy the equation.
 p – 5 = 5                          (0, 10, 5 – 5)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

p – 5 = 5
p = 10 is a solution for this equation because for p = 10,
p–5 = 10–5
= 5
∴Equation satisfied
p = 0 is not a solution for this equation because for p = 0,
p–5 =0–5
= -5 and not 5
p = 5 is not a solution for this equation because for p = 5
p–5 =5–5
= 0 and not 5
p = -5 is not a solution for this equation because for p = -5,
p–5 = -5–5
= - 10 and not 5

Thanks